This is one not sure how to do
Solve the system of equations, give an ordered pair
x+3y=3 (1)
x=5-3y (2) Im not sure what the numbers in parentheses are for

Answer by Alan3354(69443)

Solve the system of equations, give an ordered pair
x+3y=3 (1)
x=5-3y (2)
Im not sure what the numbers in parentheses are for
-----------
I think they're just equation numbers
x+3y=3 (1)
x=5-3y (2) --> x+3y = 5
x+3y cannot equal both 3 and 5. These are inconsistent, no solution.
If you graph them, they're parallel lines.